How efficient is a slide when the following occurs. The slide is 30m high and a little boy pushes off and slides down the slide with an initial velocity of 5.0 km/h and reaches the bottom while travelling 20 km/h.
1
Expert's answer
2020-06-09T13:19:50-0400

Calculate the theoretical value of final velocity applying conservation of energy principlr:


"mgh+\frac{mv_1^2}{2}=\frac{mv_\text{theor}^2}{2},\\
\space\\
v_\text{theor}=\sqrt{2gh+v_1^2}=24.8\text{ m/s}."

Since the real final velocity was 20 m, the efficiency is


"\eta=\frac{v_\text{real}}{v_\text{theor}}100\%=\frac{20}{24.8}100\%=81\%."